to convert her 99.9 F temperature to the equivalent reading on the celsius scale, the following calculation must be performed
1
Expert's answer
2020-12-14T14:50:59-0500

To  convert Fahrenheit to Celsius use the formula:

TC= (TF -32)*5/9=(99.9-32)*5/9= 37.72 °C


Answer: 37.72 °C